a seller of boiled eggs faces a market price,P=ksh15 average total cost of ptoduction,ATC=10 and output,q=150,what is its profit level?


1
Expert's answer
2022-01-24T08:37:46-0500

Profit level;

"=(P-ATC)\times Q\\=(15-10)\times 150\\=750"